### v1

@@ 3.1 Convergence analysis
$\|F(z_{k}^{t})\|\leqslant$ $~{}\big{(}\sqrt{\frac{12}{11}}+\frac{72}{11}\big{)}M_{k}^{t}\frac{1}{64^{k}}\|z_{0}-z^{*}\|^{2}+(\frac{2}{7}+1)\sqrt{\frac{12}{11}}M_{k}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|^{2}$
$~{}+\big{(}\sqrt{\frac{12}{11}}+1\big{)}4^{k-1}\sigma_{1}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|+\sqrt{\frac{12}{11}}\frac{1}{8^{k}}\|DF(z_{0})\|\|z_{0}-z^{*}\|$
$~{}+\sigma_{1}^{t}\|z_{0}-z^{*}\|+\frac{5\sigma_{1}^{t}}{4}\|z_{0}-z^{*}\|$
$\leqslant$ $~{}8M_{k}^{t}\frac{1}{64^{k}}\|z_{0}-z^{*}\|^{2}+2M_{k}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|^{2}+\frac{41}{16}\sigma_{1}^{t}\|z_{0}-z^{*}\|$
$~{}+\sqrt{\frac{12}{11}}\frac{1}{8^{k}}\|DF(z_{0})\|\|z_{0}-z^{*}\|,$
where the second inequality is by $\sqrt{\frac{12}{11}}+\frac{72}{11}\leqslant 8,\big{(}\frac{2}{7}+1\big{)}\sqrt{\frac{12}{11}}\leqslant 2$ and $\big{(}\sqrt{\frac{12}{11}}+1\big{)}4^{k-1}\frac{1}{8^{k}}\leqslant\big{(}\sqrt{\frac{12}{11}}+1\big{)}\frac{1}{8}\leqslant\frac{5}{16}$. By the choice of $K_{t}$ and $\sigma_{k}^{t}$, we have $\|F(z_{K_{t}}^{t})\|\leqslant\epsilon$. The proof is then completed by Step 2(e) in Algorithm 2. ∎
@@
Theorem 2.
Suppose that Assumption 1 holds. Let $\{z_{t}\}$ be a sequence generated by Algorithm 2. Set
$\sigma_{k}^{t}=\frac{\epsilon}{41D_{t}}4^{k},N_{k}^{t}=\left\lceil\big{(}\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\sigma_{k}^{t}}\big{)}^{\frac{2}{3}}\right\rceil,$
Then, we have $D_{t}\leqslant 4\|z_{0}-z^{*}\|$ for any $t\leqslant T(\epsilon)$ and
@@
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1).
Proof.
@@
Firstly, by Lemma 6, if $D_{t}\geqslant\|z_{0}-z^{*}\|$, we have $\|F(z_{t+1})\|\leqslant\epsilon$. By Step 3 in Algorithm 2, we have $D_{t}=4^{t}D_{0}$. The proof is then completed by using the definition of $T(\epsilon)$. ∎
Remark 2.
@@
Note that in the $t$-th iteration, simlar to the proof in Lemma 3, for any $k\geqslant 1$, we have $H_{0}^{t}=M_{0}^{t}=M_{0}$, $H_{k-1}^{t}\leqslant H_{k}^{t}\leqslant 2\rho$, and $M_{k-1}^{t}\leqslant M_{k}^{t}\leqslant 2\rho$. Therefore, at iteration $t$, the number of line searches for $M_{k}^{t}$ and $H_{k}^{t}$ is upper bounded by $2\log_{2}\frac{2\rho}{M_{0}}$. Theorem 2 implies that
$~{}\sum_{t=1}^{T(\epsilon)}\sum_{k=1}^{K_{t}}N_{k}^{t}$
$\leqslant$ $~{}\sum_{t=1}^{T(\epsilon)}K_{t}+\sum_{t=1}^{T(\epsilon)}\sum_{k=1}^{K_{t}}\big{(}\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\frac{\epsilon}{41D_{t}}4^{k}}\big{)}^{\frac{2}{3}}$
@@
$\leqslant$ $~{}T(\epsilon)\left\lceil\max\left\{\log_{8}\frac{256\rho\|z_{0}-z^{*}\|^{2}}{\epsilon},\log_{8}\frac{16\sqrt{12/11}\|DF(z_{0})\|\|z_{0}-z^{*}\|}{\epsilon}\right\}\right\rceil$
$~{}+T(\epsilon)\frac{64}{7}\big{(}\frac{43296\sqrt{3}\rho\|z_{0}-z^{*}\|^{2}}{\epsilon}\big{)}^{\frac{2}{3}},$
@@
where the second inequality is by $H_{k}^{t}\leqslant 2\rho$, $M_{k}^{t}\leqslant 2\rho$ and $D_{t}\leqslant 4\|z_{0}-z^{*}\|$. Then, the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the gradient norm is upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{2}}{\epsilon})^{\frac{2}{3}}$.
@@
Note that all the existing algorithms with iteration complexity results in terms of the gradient norm [25, 34, 40] made an additional assumption that the gradient is Lipschitz continuous. However, our method requires only a mild assumption that the Hessian is Lipschitz continuous. Although the homotopy-continuation cubic regularized Newton algorithmin in [25] has the best iteration complexity results in terms of the gradient norm, additional assumptions of error bound conditions are needed. For problems without assuming such error bound conditions, we improve the complexity in [34] for the convex-concave setting by a logarithmic factor. Moreover, our method achieves better iteration complexity compared to the adaptive second-order optimistic parameter free method in [40], which has an iteration complexity of $\mathcal{O}(\epsilon^{-1})$ in terms of the gradient norm.
@@ 4 Numerical experiment
In this section, we compare the numerical performance of the proposed LF-CR algorithm and the FF-CR algorithm with the extragradient (EG) method [18] and the Newton-MinMax algorithm [47] on the synthetic saddle point problem. All methods are implemented using MATLAB R2017b on a laptop with Intel Core i5 2.8GHz and 4GB memory.
(a) n=50, $\rho$=10
(b) n=50, $\rho$=50
Figure 1: Numerical results of the LF-CR algorithm, the FF-CR algorithm, the EG algorithm and the Newton-MinMax algorithm when $\rho\in\{10,50\}$.
We consider the following convex-concave minimax optimization problem [47]:
$\min_{x\in\mathbb{R}^{n}}\max_{y\in\mathbb{R}^{n}}\ f(x,y)=\tfrac{\rho}{6}\|x\|^{3}+y^{\top}(Ax-b),$ (62)
@@
where $\rho>0$, the entries of $b\in\mathbb{R}^{n}$ are generated independently from $[-1,1]$ and $A\in\mathbb{R}^{n\times n}$ is a invertible matrix. It can be verified that the function $f$ is $\rho$-Hessian Lipschitz, and admits a global saddle point $z^{\star}=(x^{\star},y^{\star})$ with $x^{\star}=A^{-1}b$ and $y^{\star}=-\frac{\rho}{2}\|x^{\star}\|(A^{\top})^{-1}x^{\star}$.
Experimental setup. In our experiment, the problem parameters are chosen as $n=50,\rho\in\{10,50\},A=I_{n}.$ Let $H_{0}$ be the initial estimate of Hessian Lipschize constant $\rho$ and $D_{0}$ be the initial estimate of $\|z_{0}-z^{*}\|$. The initial point of all algorithms is chosen as $z_{0}=z^{*}+0.1c$, where each component of $c\in\mathbb{R}^{2n\times 1}$ is random generated independently from $[-1,1]$.
For the EG algorithm, the stepsize is set to $0.01$ and $0.001$ when $\rho=10$ and $50$, respectively. For the LF-CR algorithm and the FF-CR algorithm, we set $H_{0}=1$, $D_{0}$=$1$. For fair comparison, we use the same solver to solve the subproblems of different algorithms. We use the gradient norm as the performance indicator of the algorithm.
Results. Figure 1 shows the gradient norm descent curves of the four algorithms. Obviously, the second-order methods perform better than the first-order method, i.e., the EG method. Among the three second-order methods, the LF-CR algorithm and the FF-CR algorithm perform better than the Newton-MinMax algorithm, especially when the Lipschtiz constant, $\rho$, increases. The reason is that the Newton-MinMax algorithm uses an accurate global Hessian Lipschitz constant, which makes the step size selection of the algorithm more conservative, while the LF-CR algorithm only estimates the Hessian Lipschitz constant locally, giving more freedom to the step size selection. Although the FF-CR algorithm is completely parameter-free, its performance is similar to that of the LF-CR algorithm, which shows its efficiency.
@@ 5 Conclusions
In this paper, we propose parameter-free second-order algorithms for unconstrained convex-concave minimax optimization problems with optimal iteration complexity. Under a mild assumption that the Hessian is Lipschitz continuous, the proposed LF-CR algorithm achieves an iteration complexity of $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{3}}{\epsilon})^{\frac{2}{3}}$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap without prior knowledge of the Lipschitz constants, and the proposed FF-CR algorithm achieves an iteration complexity of $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{2}}{\epsilon})^{\frac{2}{3}}$ in terms of the $\epsilon$-optimal solution of the gradient norm without any prior knowledge of problem parameters. To the best of our knowledge, the FF-CR algorithm is the first fully parameter-free second-order method for solving minimax optimization problems.
The only disadvantage of the proposed FF-CR algorithm is that it requires the exact optimal solutions of the subproblems. If the subproblem is allowed to be solved inexactly, whether the fully parameter-free algorithm with the optimal iteration complexity can still be obtained is worth further study.

### v2

@@ 3.1 Convergence analysis
$\|F(z_{k}^{t})\|\leqslant$ $\leavevmode\nobreak\ \bigg{(}\sqrt{\frac{12}{11}}+\frac{72}{11}\bigg{)}M_{k}^{t}\frac{1}{64^{k}}\|z_{0}-z^{*}\|^{2}+(\frac{2}{7}+1)\sqrt{\frac{12}{11}}M_{k}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|^{2}$
$\leavevmode\nobreak\ +\bigg{(}\sqrt{\frac{12}{11}}+1\bigg{)}4^{k-1}\sigma_{1}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|+\sqrt{\frac{12}{11}}\frac{1}{8^{k}}\|DF(z_{0})\|\|z_{0}-z^{*}\|$
$\leavevmode\nobreak\ +\sigma_{1}^{t}\|z_{0}-z^{*}\|+\frac{5\sigma_{1}^{t}}{4}\|z_{0}-z^{*}\|$
$\leqslant$ $\leavevmode\nobreak\ 8M_{k}^{t}\frac{1}{64^{k}}\|z_{0}-z^{*}\|^{2}+2M_{k}^{t}\frac{1}{8^{k}}\|z_{0}-z^{*}\|^{2}+\frac{41}{16}\sigma_{1}^{t}\|z_{0}-z^{*}\|$
$\leavevmode\nobreak\ +\sqrt{\frac{12}{11}}\frac{1}{8^{k}}\|DF(z_{0})\|\|z_{0}-z^{*}\|,$
where the second inequality is by $\sqrt{\frac{12}{11}}+\frac{72}{11}\leqslant 8,\left(\frac{2}{7}+1\right)\sqrt{\frac{12}{11}}\leqslant 2$ and $\left(\sqrt{\frac{12}{11}}+1\right)4^{k-1}\frac{1}{8^{k}}\leqslant\left(\sqrt{\frac{12}{11}}+1\right)\frac{1}{8}\leqslant\frac{5}{16}$. By the choice of $K_{t}$ and $\sigma_{k}^{t}$, we have $\|F(z_{K_{t}}^{t})\|\leqslant\epsilon$.
Combining (3.1) with (63) and (64), it follows that for any $k\geqslant 1$
$\|z_{k}^{t}-z^{*}\|\leqslant\frac{1}{8}\|z_{0}-z^{*}\|+\frac{2}{7}\|z_{0}-z^{*}\|+2\|z_{0}-z^{*}\|=\frac{135}{56}\|z_{0}-z^{*}\|.$
The proof is then completed by Step 3(e) in Algorithm 2.
@@
Theorem 2
Suppose that Assumptions 1 and 3 hold. Let $\{z_{t}\}$ be a sequence generated by Algorithm 2. Set
$\sigma_{k}^{t}=\frac{\epsilon}{41D_{t}}4^{k},N_{k}^{t}=\left\lceil\left(\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\sigma_{k}^{t}}\right)^{2/3}\right\rceil,$
$K_{t}=\left\lceil\max\left\{\log_{64}\frac{32M_{k}^{t}D_{t}^{2}}{\epsilon},\log_{8}\frac{8M_{k}^{t}D_{t}^{2}}{\epsilon},\log_{8}\frac{4\sqrt{\frac{12}{11}}\|DF(z_{0})\|D_{t}}{\epsilon}\right\}\right\rceil.$
Then, we have $D_{t}\leqslant 4\sqrt{\frac{12}{11}}\|z_{0}-z^{*}\|$ for any $t\leqslant T(\epsilon)$ and
@@
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1) and $D_{0}\leqslant\sqrt{\frac{12}{11}}\|z_{0}-z^{*}\|$.
@@
Proof  By Step 2 in Algorithm 2, we initialize $D_{0}$ by a backtracking routine. According to a similar proof as in (55), we have $D_{0}=\|z_{0,i}-z_{0}\|\leqslant\sqrt{\frac{12}{11}}\|z_{0}-z^{*}\|$. Notice that by Lemma 7, if $D_{t}\geqslant\|z_{0}-z^{*}\|$, we have $\|F(z_{t+1})\|\leqslant\epsilon$. The proof is then completed by using the definition of $T(\epsilon)$ and Step 4 in Algorithm 2 that $D_{t}=4^{t}D_{0}$.
Remark 2
@@
Note that in the $t$-th iteration, similar to the proof in Lemma 4, for any $k\geqslant 1$, we have $H_{0}^{t}=M_{0}^{t}=M_{0}$, $H_{k-1}^{t}\leqslant H_{k}^{t}\leqslant 2\rho$, and $M_{k-1}^{t}\leqslant M_{k}^{t}\leqslant 2\rho$. Therefore, on the one hand, at iteration $k$, the number of line search steps for $M_{k}^{t}$ and $H_{k}^{t}$ is upper bounded by $1+\log_{2}\frac{M_{k}^{t}}{M_{k-1}^{t}}$ and $N_{k}^{t}+\log_{2}\frac{H_{k}^{t}}{H_{k-1}^{t}}$, respectively. On the other hand, Theorem 2 implies that
$\leavevmode\nobreak\ \sum_{t=1}^{T(\epsilon)}\sum_{k=1}^{K_{t}}\left(N_{k}^{t}+\log_{2}\frac{H_{k}^{t}}{H_{k-1}^{t}}+\log_{2}\frac{M_{k}^{t}}{M_{k-1}^{t}}+1\right)$
@@
$\leqslant$ $\leavevmode\nobreak\ 2\sum_{t=1}^{T(\epsilon)}K_{t}+\sum_{t=1}^{T(\epsilon)}\sum_{k=1}^{K_{t}}\left(\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\frac{\epsilon}{41D_{t}}4^{k}}\right)^{2/3}+2\sum_{t=1}^{T(\epsilon)}\log_{2}\frac{2\rho}{M_{0}}$
$\leqslant$ $\leavevmode\nobreak\ 2\sum_{t=1}^{T(\epsilon)}K_{t}+\sum_{t=1}^{T(\epsilon)}D_{t}^{4/3}\sum_{k=1}^{K_{t}}2^{6-3k}\left(\frac{2706\sqrt{3}\rho}{\epsilon}\right)^{2/3}+2\sum_{t=1}^{T(\epsilon)}\log_{2}\frac{2\rho}{M_{0}}$
$\leqslant$ $\leavevmode\nobreak\ T(\epsilon)\left\lceil\max\left\{\log_{8}\frac{280\rho\|z_{0}-z^{*}\|^{2}}{\epsilon},\log_{8}\frac{18\|DF(z_{0})\|\|z_{0}-z^{*}\|}{\epsilon}\right\}\right\rceil$
$\leavevmode\nobreak\ +\frac{512}{7}\left(\frac{2706\sqrt{3}\rho\|z_{0}-z^{*}\|^{2}}{\epsilon}\right)^{2/3}+2T(\epsilon)\log_{2}\frac{2\rho}{M_{0}},$ (66)
@@
where the last inequality is by $\sum_{t=1}^{T(\epsilon)}D_{t}^{4/3}=\sum_{t=1}^{T(\epsilon)}D_{0}^{4/3}4^{4t/3}\leqslant 8\|z_{0}-z^{*}\|^{4/3}$ and $\sum_{k=1}^{K_{t}}2^{6-3k}\leqslant\frac{64}{7}.$ Then, the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the gradient norm is upper bounded by $\mathcal{O}\big{(}\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3}\big{)}$.
Remark 3
Compared with the restricted primal-dual gap termination criterion, the gradient norm is a more general termination criterion in optimization problems and is easy to compute in practice (Yoon and Ryu, 2021; Chen and Luo, 2024). This is particularly important for parameter-free optimization algorithms.
[…]
$\textsc{gap}(x_{t+1},y_{t+1};D)\leqslant\max_{x\in\mathbb{B}_{D}(x^{\star}),y\in\mathbb{B}_{D}(y^{\star})}\langle z_{t+1}-z,F(\hat{z})\rangle\leqslant 3D\|F(z_{t+1})\|\leqslant 3D\epsilon.$
Thus, by further combining the bound (2), $z_{t+1}$ becomes an $\mathcal{O}(\epsilon D)$-optimal solution with respect to the restricted primal-dual gap after no more than $\mathcal{O}\left(\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3}\right)$ additional iterations of the FF-CR algorithm. This means that the gradient norm termination criterion is stronger than the primal-dual gap in this sense.
@@
Note that although the homotopy-continuation cubic regularized Newton algorithm in (Huang et al., 2022) has the best iteration complexity results in terms of the gradient norm, additional assumptions of error bound conditions are needed. For problems without assuming such error bound conditions, we improve the complexity in (Ostroukhov et al., 2020) for the convex-concave setting by a logarithmic factor. Moreover, our method achieves better iteration complexity compared to the adaptive second-order optimistic method in (Jiang et al., 2024), which has an iteration complexity of $\mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)$ in terms of the gradient norm.
@@ 4 Numerical experiment
In this section, we compare the numerical performance of the proposed LF-CR algorithm and the FF-CR algorithm with the extragradient (EG) method (Korpelevich, 1976), the Newton-MinMax method (Lin et al., 2022), the homotopy inexact proximal-Newton extragradient (HIPNEX) algorithm (Alves et al., 2024), the optimistic second-order method (OSOM) with line search (Jiang and Mokhtari, 2022) and the parameter-free variant of the adaptive second-order optimistic method (ASOM) (Jiang et al., 2024) in solving a synthetic minimax optimization problem and an AUC maximization problem. All methods are implemented using MATLAB R2017b on a laptop with Intel Core i5 2.8GHz and 4GB memory.
4.1 A synthetic minimax optimization problem
We consider the following convex-concave minimax optimization problem (Lin et al., 2022):
$\min_{x\in\mathbb{R}^{n}}\max_{y\in\mathbb{R}^{n}}\ f(x,y)=\tfrac{\rho}{6}\|x\|^{3}+\langle y,Ax-b\rangle,$ (67)
where $\rho>0$, the entries of $b\in\mathbb{R}^{n}$ are generated independently from [1,1][-1,1] and $A\in\mathbb{R}^{n\times n}$ is given by
$A=\begin{bmatrix}1&-1&&&\\ &1&-1&&\\ &&\ddots&\ddots&\\ &&&1&-1\\ &&&&1\end{bmatrix}$
@@ 4.1 A synthetic minimax optimization problem
It can be verified that the function $f(x,y)$ is $\rho$-Hessian Lipschitz continuous, and admits a global saddle point $z^{\star}=(x^{\star},y^{\star})$ with $x^{\star}=A^{-1}b$ and $y^{\star}=-\frac{\rho}{2}\|x^{\star}\|(A^{\top})^{-1}x^{\star}$.
In our experiment, the problem parameters are chosen as $n\in\{50,100\},\rho=\frac{1}{15n}$. Let $H_{0}$ be the initial estimate of Hessian Lipschitz constant $\rho$ and $D_{0}$ be the initial estimate of $\|z_{0}-z^{*}\|$. For the LF-CR algorithm and the FF-CR algorithm, after selecting the initial point $z_{0}$, another random point $\tilde{z}_{0}$ close to the initial point is generated. Then, $H_{0}$ and $D_{0}$ are generated using the formulas in the two algorithms, and each cubic subproblem is calculated until a high accuracy is achieved. For the HIPNEX algorithm, we set $\hat{\sigma}=0$, $\theta=0.3$, and other hyperparameters are determined by formulas $(24)$ and $(25)$ in Alves et al. (2024). For the OSOM algorithm, we set the initial step size $\sigma_{0}=1$ and the line search hyperparameters $\alpha=0.5$ and $\beta=0.3$. For the ASOM algorithm, we set $\lambda_{0}=H_{0}$. We run each algorithm until $\|F(z)\|\leqslant 1e-4.$
(a) nn=50
(b) nn=100
Figure 1: Numerical results of the tested algorithms for solving synthetic minimax problem
Figure 1 shows the gradient norm descent curves of the seven algorithms. Obviously, the second-order methods perform better than the first-order method, i.e., the EG method. Among the second-order methods, the LF-CR algorithm and the FF-CR algorithm perform slightly better than the other methods.
4.2 AUC maximization problem
We consider the following minimax formulation for the area under the receiver operating characteristic curve (AUC) maximization which aims to find a classifier $\theta\in\mathbb{R}^{n}$ that maximizes AUC score on a set of samples $\{(a_{i},b_{i})\}_{i=1}^{M}$, where $a_{i}\in\mathbb{R}^{n}$ and $b_{i}\in\{1,-1\}$ (Lin et al., 2022; Ying et al., 2016; Shen et al., 2018):
$\min_{x=(\theta,u,v)}\max_{y}$ $\frac{1-p}{M}\left\{\sum_{i=1}^{M}(\langle\theta,a_{i}\rangle-u)^{2}\mathbb{I}_{[b_{i}=1]}\right\}+\frac{p}{M}\left\{\sum_{i=1}^{M}(\langle\theta,a_{i}\rangle-v)^{2}\mathbb{I}_{[b_{i}=-1]}\right\}$
$+\frac{2(1+y)}{M}\left\{\sum_{i=1}^{M}\langle\theta,a_{i}\rangle\bigg{(}p\mathbb{I}_{[b_{i}=-1]}-(1-p)\mathbb{I}_{[b_{i}=1]}\bigg{)}\right\}+\frac{\rho}{6}\|x\|^{3}-p(1-p)y^{2},$
where $\rho>0$ is a scalar, $u,v\in\mathbb{R}$ are auxiliary variables, $\mathbb{I}_{[\cdot]}$ is the indicator function, and $p$ is the proportion of samples with positive label.
In our experiment, we choose $\rho=\frac{1}{M}$, which is the same as in (Lin et al., 2022), and use two imbalanced binary classification datasets “a9a” $(n=126,M=32561)$, “w8a” $(n=300,M=49749)$ from LIBSVM datasets. For the LF-CR algorithm and the FF-CR algorithm, we generate $H_{0}$, $M_{0}$ and $D_{0}$ using the formulas in both algorithms. For the HIPNEX algorithm, we set $\hat{\sigma}=0$ , $\theta=0.2$, and other hyper-parameters are determined by the formulas $(24)$ and $(25)$ in Alves et al. (2024). For the OSOM algorithm, we set the initial stepsize $\sigma_{0}=1$, the line-search hyperparameter $\alpha=0.5$ and $\beta=0.5.$ For the ASOM algorithm, we set $\lambda_{0}=H_{0}$. We run each algorithm until $\|F(z)\|\leqslant 1e-10.$
Figure 2 shows the numerical performance of the seven tested algorithms when solving the AUC maximization problem, where the horizontal axis represents the number of iterations and the vertical axis represents the gradient norm of $f(x,y)$. The results show that the proposed FF-CR algorithm slightly outperforms the other algorithms.
(a) a9a
(b) w8a
Figure 2: Numerical results of the tested algorithms for solving AUC maximization problem
@@ 5 Conclusions
In this paper, we propose parameter-free second-order algorithms for unconstrained convex-concave minimax optimization problems. Under a mild assumption that the Hessian is locally Lipschitz continuous, the proposed LF-CR algorithm achieves the iteration complexity of $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ (resp. $\mathcal{O}(\rho\|z_{0}-z^{*}\|^{2}\epsilon^{-1})$) in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm) without prior knowledge of the Lipschitz constants. And the proposed FF-CR algorithm achieves the iteration complexity of $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3})$ in terms of the $\epsilon$-optimal solution of the gradient norm without any prior knowledge of problem parameters. To the best of our knowledge, the proposed FF-CR algorithm is a completely parameter-free second-order algorithm, and its iteration complexity is currently the best in terms of $\epsilon$ under the termination criterion of the gradient norm.
The only disadvantage of the proposed FF-CR algorithm is that it requires the exact optimal solutions of the subproblems. If the subproblems are allowed to be solved inexactly, whether it is still possible to obtain the current iteration complexity of a completely parameter-free algorithm is worth further study.